OpenStudy (anonymous):

find all the solutions to the interval [0,2pi) of the equation cos^22x=1/2

OpenStudy (anonymous):

cos^2(2x)=1/2 cos2x=(+/-)1/4 case 1, cos 2x=1/4, 2x=1.32,4.97 x=0.66,2.49 case 2, cos 2x=-1/4, 2x=1.82,3.45 x=0.91,1.73

OpenStudy (anonymous):

dont know if there are some more solution. I haven't done these for long